1.Various advanced texts in linear algebra prove the following determinant criterion for rank:

The rank of a matrix A is r if and only if A has some r × r sub matrix with a nonzero determinant, and all square sub matrices of larger size have determinant zero.

(A sub matrix of A is any matrix obtained by deleting rows or columns of A. The matrix A itself is also considered to be a sub matrix of A.) Use this criterion to find the rank of the matrix.

1 4 0
3 12 -1
rank (A) =


rank (A) =
1
Expert's answer
2020-01-10T09:44:53-0500

"\begin{bmatrix}
 1 & 4 & 0\\
 3 & 12 & -1
\end{bmatrix}" is the given matrix. Performing the operation "R_2 \gets R_2-3R_1" we get;

"\begin{bmatrix}
 1 & 4 & 0\\
 0 & 0 & -1
\end{bmatrix}" . Now, the sub matrix "\begin{bmatrix}
 4 & 0 \\
 0 & -1
\end{bmatrix}" of order "2\Chi2" has a non-zero determinant.

Also, this is the largest sub matrix with non-zero determinant.

Thus, rank of the given matrix is 2.